OpenStudy (anonymous):

The ratio of children to adults at a baseball game is 2 to 5. If there are 7,200 children at the game, which of the following could not be used to determine the number of adults, a? 2/5=7200/a 2/a=5/7200 7200/2=a/5 a/7200=5/2

OpenStudy (anonymous):

the first one cannot be used

OpenStudy (anonymous):

no - i think that one can be used if we cross multipky 2a = 5* 7200 a = 5*7200 / 2

OpenStudy (anonymous):

which would give you like 18,000 adults

OpenStudy (anonymous):

yup adults are in majority

OpenStudy (anonymous):

i read the question wrong...i thought it was 7200 total people at the game

OpenStudy (anonymous):

for every 2 children there are 5 adults

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so then yea...it's the 2nd equation that can't be used

OpenStudy (anonymous):

yup - 2nd one